OpenStudy (anonymous):

what is 15/2 divided by 69/8?

hero (hero):

You are tasked with simplifying \(\dfrac{15}{2} \div \dfrac{69}{8}\) right?

OpenStudy (anonymous):

I believe so yes

hero (hero):

Are you familiar with the reciprocal rule of dividing fraction where if have \(\dfrac{a}{c} \div \dfrac{d}{b}\) then it is the same as \(\dfrac{a}{c} \times \dfrac{b}{d}\)

OpenStudy (anonymous):

yes

hero (hero):

So according to this rule, \(\dfrac{15}{2} \div \dfrac{69}{8} = \dfrac{15}{2} \times \dfrac{8}{69}\) right?

OpenStudy (anonymous):

yes so would you just divide 15 by 8 and 2 by 69 then simplify?

hero (hero):

At every step, you want to simplify first before attempting to perform any operation. There's another rule you should be aware of and that is cross cancellation. But just to make the rule clearer, if we have \(\dfrac{a}{c} \times \dfrac{b}{d}\), we can write it as one fraction \(\dfrac{ab}{cd}\) Furthermore, by the associative property, the values being multiplied are interchangable, therefore if we want we can write \(\dfrac{ab}{cd} \) as \(\dfrac{ba}{dc}\)

hero (hero):

In your case, to help simplify things easier, we can write \(\dfrac{15}{2} \times \dfrac{8}{69} = \dfrac{(15)(8)}{(2)(69)}\) And then swap the values of the numerator to get \(\dfrac{(8)(15)}{(2)(69)}\) Do you follow any of this so far?

OpenStudy (anonymous):

then you just divide?

OpenStudy (anonymous):

@Hero then you just divide?

hero (hero):

Well again, we want to simplify if possible before dividing and if the numerator is smaller than the denominator, then it is appropriate to simplify as much as possible and then leave the the result in fraction form.

hero (hero):

Now, swapping the values has a valid purpose because we can now write \(\dfrac{(8)(15)}{(2)(69)}\) as \(\dfrac{8}{2} \times \dfrac{15}{69}\)

hero (hero):

And in this case we can divide 8/2, but we are limited to simplifying 15/69. Do you see this?

OpenStudy (anonymous):

yeah cuz 15/69 is a decimal and you cant have a decimal in a fraction

hero (hero):

It is not necessary to simplify it down to decimals. We want to reduce it to the point where we can just leave it in fraction form.

hero (hero):

Obviously 8/2 = 4 so what we have left is \(4 \times \dfrac{15}{69}\) Do you get that?

OpenStudy (anonymous):

I get how you got 4 but what do you do with the 15/69?

hero (hero):

We simplify it by cancelling "factors of one". Notice that \(\dfrac{15}{69} = \dfrac{(3)(5)}{(3)(23)}\) Do you get that part?

OpenStudy (anonymous):

yes

hero (hero):

A particular thing to notice here is that \(\dfrac{(3)(5)}{(3)(23)} = \dfrac{3}{3} \times \dfrac{5}{23}\)

hero (hero):

And obviously \(\frac{3}{3} = 1\)

OpenStudy (anonymous):

so \[1 \times 5/23\]

hero (hero):

But let's not forget the 4 because what remains is: \(4 \times \dfrac{5}{23}\)

OpenStudy (anonymous):

oh so would you do this: \[4/1 \times 5/23?\]

hero (hero):

Correct, so we can multiply that together since we are done simplifying. Can you compute the final result here?

OpenStudy (anonymous):

20/23?

hero (hero):

Great job. So to summarize, the process of multiplying and dividing two fractions, is actually the art of reducing those fractions to their most simplest forms by cancelling factors of one first before performing the appropriate operations on them. Once you have mastered this, then you'll be able to multiply or divide any fractions with confidence.